### v1

# Jet launching from the Kerr black hole magnetosphere: An electrogeodesic approach

Abstract
The launch of relativistic jets of plasma on astrophysical to cosmological scales are observed in a variety of astrophysical sources, from active galactic nuclei to X-rays binaries. While these jets can be reproduced by the general relativistic magneto-hydrodynamic (GRMHD) and particle-in-cells (GRPIC) simulations of the dynamical Kerr magnetosphere, the development of analytic models to describe the physics of the jets has remained limited. A key challenge is to analytically describe the individual trajectories of accelerated charged particles which ultimately build up the jet and emit radiation. In this work, we provide a first simple but fully analytical model of jet launching from the Kerr magnetosphere based on the motion of charged particles. To that end, we use the integrability of electrogeodesic motion in the Kerr monopole magnetosphere to study the ejection of charged particles near the poles. This enables us to derive (i) a criterion for the rotation axis to constitute a stable latitunal equilibrium position, thereby representing an idealized jet, (ii) the expression for the magnetic frame-dragging effect, and (iii) the condition for an asymptotic observer to measure blueshifted particles emanating from the black hole surroundings. Our study reveals that particles can be accelerated only in a specific region whose maximal radius depends on the spin and magnetization of the black hole. Alongside these results, we provide a detailed review of the construction of test magnetospheres from (explicit and hidden) symmetries of the Kerr geometry and the condition for the separability of the electrogeodesic motion in a test magnetosphere, which serves as a basis for the model we study in this work.
Introduction
Relativistic jets of plasma are observed in a variety of astrophysical systems, from active galactic nuclei (AGN) containing a supermassive black hole (SMBH) in its centter to stellar mass black hole involved in X-rays binaries [1, 2]. They can extend up to thousand of light years, therefore playing a major role in galaxy evolution by redistributing matter on large scales [3]. These jets are believed to be powered by highly magnetized spinning black holes which accrete plasma and accelerate particles at very high energy. Yet the precise mechanisms at play to produce and sustain these powerful jets, together with a detailed understanding of their structure, still remains to be understood [4]. While catalogue of astrophysical systems powering jets have been monitored during the last decades [5, 6], the new born horizon-scale electromagnetic astronomy initiated by the Event Horizon Telescope [7, 8] now enables us to image relativistic jets of various systems with an unprecedented resolution, among which the M87 SMBH [9]. These observation coupled to a better characterization of the properties of the jet could provide a new window to further constrain the mass and spin of SMBH [10].
To a large extent, our theoretical understanding of energy extraction from highly magnetized spinning black hole relies on the seminal Blandford-Znajek (BZ) model [11]. The BZ solution describes a slowly rotating monopole magnetosphere surrounding the Kerr black hole, which is sourced by a dilute plasma. This Kerr magnetosphere successfully generates outgoing Poynting fluxes triggered by the toroidal components of the magnetic field. The BZ model relies on the force-free (FF) regime where the electromagnetic field satisfies $E\cdot B=0$ and $E^{2}<B^{2}$ everywhere. The FF approximation ensures that the plasma density $\rho$ is sufficient to source the magnetosphere but small enough to keep it in a high magnetization regime, i.e. $\sigma\sim B^{2}/\rho\gg 1$. See [12, 13, 14, 15, 16, 17] for more details on the BZ process. While the validity of the FF approximation and the asymptotic stability of the BZ solution casted some doubts in the early days [18], the progress in general relativistic (GR) magneto-hydrodynamics (MHD) simulations of dynamical black hole magnetospheres [19, 20, 21] allowed to address these questions and eventually numerically reproduce the final steady-state BZ solution and its Poynting winds [22, 23, 24, 25]. See [26] for a pedagogical review on the theoretical description of force-free magnetosphere.
Nevertheless, the BZ model and its associated GRMHD simulations suffer from several limitations to properly characterize the jet launching mechanism. First, the BZ solution is perturbative in the spin and thus cannot account for high spinning black hole. So far, no non-perturbative extension has been found [27, 28, 29, 30]. Second, and more importantly, the GRMHD simulations rely on a fluid description of the plasma which is inherently incomplete. In particular, it fails to track the motion of individual charged particles in the magnetosphere which requires a kinetic approach. Switching from the fluid-based to a particle-based description of the jet is crucial to make progress. Moreover, the FF condition forbids the formation of pairs and electric gaps which are believed to play an important role in the jet dynamics. Even more importantly, the FF condition prevents charged particle to locally accelerate, which happens only when $E\cdot B\neq 0$, i.e. in the presence of non-ideal electric fields. The developments of GR/Particle-In-Cells (PIC) simulations have played a major role in addressing these difficulties [31, 32, 33]. These simulations, based on the ZELTRON code, allow us to follow the individual particles trajectories of the plasma evolving in a dynamical magnetosphere where the force free condition can be violated. While they could also reproduce jet launching from Kerr immersed in a asymptotic uniform field [34], they also revealed new effects such as the build up of current sheet at the equator where magnetic reconnection appears to play a crucial role to eject energy from the black hole. See [35, 36, 37, 38, 39, 40, 41] for further results along this line. Besides their qualitative results, these simulations underline the crucial need to properly characterize the motion of individual charged particles (as opposed to fluid) in the black hole magnetosphere to properly account of the complicated mechanisms at play in the accretion/ejection paradigm.
The important progresses in the development of the GR/PIC simulations of the black hole magnetosphere dramatically contrast with the analytic investigations, which have remained scarce [42, 43]. Indeed, the inherent complexity in solving the electrogeodesic motion around magnetized black hole has limited most studies to a mixture of analytical and numerical treatments [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. This prevents us to obtain close analytic predictions for the structure of the jet, its opening angle, the size of the acceleration and escape zones or the azimuthal velocity of charged particles in the different regions of the jet, which should play a key role to describe the emitted radiation. In contrast, the neutral geodesic motion in the Kerr geometry is well known to be fully integrable analytically [54]. The existence of a hidden symmetry, leading to the famous Carter constant, allows one to fully separate the radial and polar motion, which in turn, allows one to separate the time and azimuthal motion. In general, this two step process breaks down when a test-field magnetosphere is turned on, such that the electrogeodesic equations are no longer separable. This is typically the case of the well-known and extensively studied Wald solution [66], which describes an asymptotic uniform magnetic field around a Kerr black hole. However, the situation is not completely hopeless.
Indeed, as initially shown in the seminal work of Carter [54], the separability of the geodesic motion on the Kerr black hole extends to the electrogeodesic motion on the electrically and magnetically charged Kerr-Newman (KN) black hole, i.e. the electrovacuum solution to the Einstein-Maxwell equations extending the Kerr solution and parametrized by four constants: the mass $M$, the spin $a$, the electric charge $\mathcal{Q}$} and the magnetic charge $\mathcal{P}$}. To our knowledge, the only systematic effort to extend Carter’s analytic result and classify the electrogeodesic motion in the KN geometry was presented much later by Hackman and Xu in [55]. See [56, 57, 59, 58] for related but more specific investigations. At first, it would appear natural to use these elegant results to study the jet formation in the KN magnetosphere. However, the presence of an electric charge $\mathcal{Q}$} makes the model not so realistic, since the surrounding plasma would rapidly discharge the black hole. As for the magnetic charge, this limitation does not exist given that no elementary particles carrying magnetic charge are known to exist. Yet, the magnetic charge $\mathcal{P}$} of the black hole is expected to be negligible compared to its mass, i.e. ${\mathcal{P}}/M\ll 1$, so that gravity remains the dominant effect. This more realistic regime, i.e. ${\mathcal{Q}}=0$ and ${\mathcal{P}}/M\ll 1$, corresponds to the test field approximation of the magnetic KN geometry. Interestingly, taking this limit reduces the problem to the magnetically charged, vacuum magnetosphere on the Kerr black hole which, although well known, has not received much attention in the literature. Yet, it appears to be the simplest analytic starting point to model the a rotating black hole vacuum magnetosphere. At this level, a remarkable observation is that the separability of the electrogeodesic motion in the KN black hole remains valid in the test field approximation. This allows one to obtain closed form solutions to the equations of motion of charged particles. In this work, we use this path to analytically study the acceleration and the ejection of test charged particles around a magnetized Kerr black hole. Surprisingly, even in this simple vacuum magnetosphere, the exact electrogeodesic solutions reveal a rich behavior and allows one to identify a subset of trajectories that build up a jet of accelerated particles. Let us summarize the new qualitative results for the physics of the jet emerging from this study:
  • We present the analytic expression for the $4$-velocity $u^{\mu}\partial_{\mu}$ [Eq. (4.6) below] and 4-acceleration $a^{\mu}\partial_{\mu}$ [Eq. (4.7)] of charged particles in the whole black hole exterior, which allows us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, such that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of jet of charged particles preferentially around the poles. For negative charged particles, $\kappa<0$, the acceleration is outwards and the particles are ejected if ${\mathcal{P}}>0$.
  • To understand the structure of the ejection at the poles, and thus of the jet, we further analyze the stable and unstable equilibrium positions of the polar motion depending on the energy and angular momentum $(\epsilon,\ell)$ of the charged particle. The presence of the magnetosphere dramatically changes the properties of the polar potential. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$ where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the Northern rotation axis $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we identify a new condition [Eq. (4.22) below] for the Northern rotation axis to be a stable latitudinal position, namely the energy of the particle must be either smaller or greater than a threshold depending on $\beta$. Instead, trajectories violating this condition admit a stable latitudinal position at some $\theta_{\ast}\in\;]0,\pi/2[$.
  • We also study the fate of the frame-dragging effect in the Kerr monopole. Using the expression of the azimuthal angular velocity [Eq. (4.27) below], we show that the magnetic monopole introduces a new magnetic frame-dragging which completely dominates the gravitational frame-dragging. Indeed, we find that the former decays as $1/r^{2}$ while the later decays as $1/r^{3}$, showing that the new magnetic frame-dragging impacts charged particles on larger scales. Furthermore, in contrast to the gravitational frame-dragging which forces the particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their charge, i.e. on $\kappa$.
  • We further characterize the condition under which charged particles ejected from the magnetosphere reach an asymptotic observer with a higher energy. We evaluate the redshift experienced by the particle and find the condition for the particle to be blueshifted at large distance. Interestingly, this condition reveals the existence of a structure within the magnetosphere. It manifests through the existence of a radius which encodes the maximal extent of the region in which emitted particles escaping to infinity will be observed blueshifted.
These different points, which rely on explicit analytic expressions, reveal a fine structure of the ejection mechanism of charged particles in the Kerr magnetosphere. It provides a first fully analytical study of the jet launching mechanism from the Kerr magnetosphere based on the exact solutions to the electrogeodesic motion.
Since this investigation is made possible thanks to the separability of the electrogeodesic motion in this specific model of the Kerr magnetosphere, it is interesting to understand the fundamental reasons why this elegant integrability occurs. While this is well known for the Kerr-Newman geometry, it is interesting to revisit this question from the point of view of the test field approximation. Indeed, other exact solutions such as the Wald solution which has been extensively studied fails to have separable electrogeodesic motion. To make this work as self-contained as possible, we present, prior to our study of charged particle ejection, a review of (i) the construction of exact solutions of the Maxwell equations in the Kerr geometry using the Killing symmetries and (ii) a detailed derivation of the geometrical conditions for preserving the separability, i.e. the Carter constant, for the electrogeodesic motion in a given test field magnetosphere. We also include a detailed derivation of the separable electrogeodesic equations in the Kerr magnetic monopole magnetosphere. This review allows us to present, besides the well known Wald construction, other results which have not received the same focus, such as the construction of exact vacuum solutions of Maxwell equations using Killing-Yano tensors known as the Penrose currents. We also use this opportunity to introduce new results, such as a new construction to build an exact non-vacuum solution of the Maxwell equation using Killing vectors, leading to a new exact Maxwell solution on the Kerr geometry [Eqs. (2.42)–(2.43) below]. In the end, this review provides the necessary technical tools to appreciate the fundamental reasons ensuring the separability of the electrogeodesic motion in the Kerr monopole magnetosphere, which descends from the properties of the Killing-Yano tensor of the Kerr geometry.
This work is organized as follows. In Section 2, we present the detailed review of the construction of exact test magnetospheres models from the explicit and hidden Killing symmetries, i.e. using the Killing vectors and the Killing-Yano tensors and apply them to the Kerr geometry. We present an alternative to the Wald construction in Section 2.1.3. Section 3 is devoted to the review on the separability of the electrogeodesic motion. We first review how the standard conditions for the geodesics are modified for the electrogeodesic case. Then, we derive the geometric condition for preserving the Carter constant. The detailed demonstration of the separability in the Kerr monopole magnetosphere is rederived in Section 3.4, together with the integral solutions for the electrogeodesic. In Section 4, we apply this result to study the ejection of charged particles from the Kerr magnetosphere and the structure of the jet. Finally, Section 5 provides a discussion of our results and the open directions to explore.
Kerr magnetospheres from symmetries
In this section, we review the key constructions allowing one to obtain exact vacuum solutions of Maxwell equations using the symmetries of the spacetime. Consider therefore the Maxwell equations in a curved spacetime which take the form
$\nabla_{\mu}F^{\mu\nu}$ $=J^{\mu}$ (2.1)
$\nabla_{\mu}F_{\nu\alpha}+\nabla_{\nu}F_{\alpha\mu}+\nabla_{\alpha}F_{\mu\nu}$ $=0$ (2.2)
$\nabla_{\mu}J^{\mu}$ $=0$ (2.3)
where the Faraday tensor $F_{\mu\nu}$ can be decomposed in term of the gauge-potential $A=A_{\mu}\textrm{d}x^{\mu}$ as
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ (2.4)
See [60] for a review of Maxwell theory in curved spacetime.
In the following, we shall review how to construct specific solutions of these equations. The focus will be on the Kerr metric which reads
$\textrm{d}s^{2}$ $=-\left(1-\frac{2Mr}{\Sigma}\right)\textrm{d}t^{2}-\frac{4aMr\sin^{2}{\theta}}{\Sigma}\textrm{d}t\textrm{d}\varphi+\frac{\Sigma}{\Delta}\textrm{d}r^{2}+\Sigma\textrm{d}\theta^{2}$
$\qquad\qquad+\left(r^{2}+a^{2}+\frac{2a^{2}Mr\sin^{2}{\theta}}{\Sigma}\right)\sin^{2}{\theta}\textrm{d}\varphi^{2}$ (2.5)
where the functions $\Sigma$ and $\Delta$ read
$\Sigma=r^{2}+a^{2}\cos^{2}{\theta}\qquad\Delta=r^{2}-2Mr+a^{2}=(r-r_{+})(r-r_{-})$ (2.6)
with $r_{\pm}$ the positions of the outer and inner black hole horizons. It will also be useful to write its inverse form which reads
$g^{tt}$ $=-\frac{1}{\Delta}\left(r^{2}+a^{2}+\frac{2a^{2}Mr\sin^{2}{\theta}}{\Sigma}\right)\qquad g^{\varphi\varphi}=\frac{1}{\Delta\sin^{2}{\theta}}\left(1-\frac{2Mr}{\Sigma}\right)$ (2.7)
$g^{t\varphi}$ $=-\frac{2Mar}{\Sigma\Delta}\qquad g^{rr}=\frac{\Delta}{\Sigma}\qquad g^{\theta\theta}=\frac{1}{\Sigma}$ (2.8)
The Kerr geometry is well known to have two commuting Killing vectors $\partial_{t}$ and $\partial_{\varphi}$. It also enjoys hidden symmetries generated by the presence of (conformal) Killing-Yano tensors. In the following, we shall review how both the Killing vectors and the Killing-Yano tensors encode the solutions of the Maxwell equations. In particular, this will allow us to illustrate the role of the Killing tower structure of the Kerr geometry in relating different known solutions of the Maxwell equations. See [62, 61, 63, 64, 65] for more details on the Killing tower construction.
Maxwell solutions from Killing vectors
Consider a Killing vector $k^{\mu}\partial_{\mu}$ defined by $\nabla_{(\mu}k_{\nu)}=0$. This definition implies that it also satisfies the following equations
$\Box k_{\mu}=-R_{\mu\nu}k^{\nu}\qquad\text{and}\qquad\nabla_{\mu}k^{\mu}=0.$ (2.9)
Writing down the source-free ($J^{\mu}=0$) Maxwell equation (2.1) in terms of the potential $A_{\mu}\textrm{d}x^{\mu}$, one finds
$\Box A_{\mu}=R_{\mu\nu}A^{\nu}.$ (2.10)
Therefore, upon working in the Lorentz gauge where $\nabla_{\mu}A^{\mu}=0$, it is direct to see that any Killing vector of a metric obeying the vacuum Einstein equation $R_{\mu\nu}=0$ provides an exact analytic solution for the source-free Maxwell equation by setting $A_{\mu}=\alpha k_{\mu}$, where $\alpha$ is a constant. Hence, one obtains a very straightforward way to relate the profile of the Maxwell field to the explicit symmetries of the vacuum geometry.
The Kerr metric being stationary and axisymmetric, it possesses two Killing vectors:
$\xi^{\mu}\partial_{\mu}=\partial_{t}\qquad\text{and}\qquad\chi^{\mu}\partial_{\mu}=\partial_{\varphi},$ (2.11)
which naturally gives rises to a two parameters family of source-free Maxwell solutions of the form
$A^{(\alpha,\beta)}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}+\beta\chi_{\mu}\textrm{d}x^{\mu}.$ (2.12)
Subsectors of this family of solutions have been first studied by Wald [^1] See [68] for an extension where the magnetosphere is misaligned w.r.t the black hole spin. in [66]. In the following, we discuss the different branches of solutions and their properties.
Before presenting the form of the electric and magnetic fields described by these solutions, it will reveal useful to introduce the Carter tetrad associated to the family of Carter observers. Each observer in this family is located at a fixed radial and polar coordinates $(r,\theta)$ and is rotating around the black hole at the angular velocity
$\Omega_{\rm C}=\frac{\textrm{d}\varphi}{\textrm{d}t}=\frac{a}{r^{2}+a^{2}}.$ (2.13)
The 4-velocity of a Carter observer is a linear combination of the two Killing vectors: $e_{0}^{\mu}=(e_{0})^{t}(\xi^{\mu}+\Omega_{\rm C}\chi^{\mu})$ and the tetrad carried by that observer is the orthonormal vector frame $(e_{\alpha})$ defined in terms of the Boyer-Lindquist coordinates by
$e_{0}$ $=\frac{r^{2}+a^{2}}{\sqrt{\Sigma\Delta}}\,\partial_{t}+\frac{a}{\sqrt{\Sigma\Delta}}\,\partial_{\varphi}$ (2.14a)
$e_{1}$ $=\sqrt{\frac{\Delta}{\Sigma}}\,\partial_{r}$ (2.14b)
$e_{2}$ $=\frac{1}{\sqrt{\Sigma}}\,\partial_{\theta}$ (2.14c)
$e_{3}$ $=\frac{a\sin\theta}{\sqrt{\Sigma}}\,\partial_{t}+\frac{1}{\sqrt{\Sigma}\sin\theta}\,\partial_{\varphi}.$ (2.14d)
The Carter observer has the distinctive feature that both the ingoing and outgoing principal null geodesics of the Kerr metric are purely radial with respect to him, i.e. their tangent vectors have no component along $e_{2}$ and $e_{3}$. This explains why the Carter observer is well adapted to the description of the Kerr spacetime (outside the black hole region). Moreover, in the asymptotic region, i.e. for $r\to+\infty$, the Carter observer coincides with an inertial observer at rest with respect to the black hole. The Carter coframe $(e^{\alpha})$ is the dual to the Carter tetrad (i.e. the unique set of 1-forms $(e^{\alpha})$ such that $(e^{\alpha})_{\mu}(e_{\beta})^{\mu}=\delta^{\alpha}_{\ \,\beta}$). It is expressed in terms of the coframe $(\textrm{d}x^{\alpha})$ associated to the Boyer-Lindquist coordinates by
$e^{0}$ $=\sqrt{\frac{\Delta}{\Sigma}}\left(\textrm{d}t-a\sin^{2}{\theta}\,\textrm{d}\varphi\right)$ (2.15a)
$e^{1}$ $=\sqrt{\frac{\Sigma}{\Delta}}\,\textrm{d}r$ (2.15b)
$e^{2}$ $=\sqrt{\Sigma}\,\textrm{d}\theta$ (2.15c)
$e^{3}$ $=\frac{\sin{\theta}}{\sqrt{\Sigma}}\left(-a\,\textrm{d}t+(r^{2}+a^{2})\,\textrm{d}\varphi\right).$ (2.15d)
We can now discuss the different types of interesting solutions belonging to the family (2.12).
Asymptotically vanishing case: the electric and magnetic monopoles
Let us first discuss the set of solution describing configurations for which $F\to 0$ for $r\to+\infty$. The first branch corresponds to $\beta=0$ in Eq. (2.12). In the following, we write
$\alpha=-\frac{{\mathcal{Q}}}{2M},$ (2.16)
such that $\mathcal{Q}$} encodes the strength of the field (actually the total electric charge, as we shall see below). Equation (2.12) then yields
$A^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{Q}}}{2M}\left(1-\frac{2Mr}{\Sigma}\right)\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.17)
Note that the term $\frac{{\mathcal{Q}}}{2M}\textrm{d}t$ is an exact differential and thus does not contribute to the electromagnetic field $F^{(\alpha,0)}$, so that a 4-potential equivalent to (2.17) is
${A^{\prime}}^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=-\frac{{\mathcal{Q}}r}{\Sigma}\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.18)
The associated Faraday tensor $F^{(\alpha,0)}=\textrm{d}A^{(\alpha,0)}=\textrm{d}{A^{\prime}}^{(\alpha,0)}$ is given terms of the Carter coframe (2.15) by
$\boxed{F^{(\alpha,0)}=2\alpha\nabla_{\mu}\xi_{\nu}\;\textrm{d}x^{\mu}\wedge\textrm{d}x^{\nu}=-\frac{{\mathcal{Q}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}\,e^{0}\wedge e^{1}+\frac{2a{\mathcal{Q}}r\cos{\theta}}{\Sigma^{2}}\,e^{2}\wedge e^{3}}$ (2.19)
A direct computation shows that this Maxwell solution has the electric charge $\mathcal{Q}$}, but no magnetic charge:
$Q_{\rm e}=\frac{1}{4\pi}\oint_{S}\ast F^{(\alpha,0)}={\mathcal{Q}}\quad\mbox{and}\quad Q_{\rm m}=\frac{1}{4\pi}\oint_{S}F^{(\alpha,0)}=0.$ (2.20)
The electric and magnetic field measured by the Carter observer are respectively $E^{(\alpha,0)\mu}=F^{(\alpha,0)\mu}_{\phantom{(\alpha,0)\mu}\,\nu}(e_{0})^{\nu}$ and $B^{(\alpha,0)\mu}=-\ast F^{(\alpha,0)\mu}_{\phantom{(\alpha,0)\mu}\,\nu}(e_{0})^{\nu}$. Given that $(e_{\alpha})$ is an orthonormal tetrad and $(e^{\alpha})$ is its dual basis, we can read them directly on (2.19):
$E^{(\alpha,0)}$ $=\frac{{\mathcal{Q}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}\,e_{1}$ (2.21)
$B^{(\alpha,0)}$ $=\frac{2a{\mathcal{Q}}r\cos{\theta}}{\Sigma^{2}}\,e_{1}.$ (2.22)
They are both purely radial and vanish asymptotically. When taking the non-rotating (Schwarzschild) limit, i.e. $a=0$, one finds that
$A^{(\alpha,0)}=\frac{{\mathcal{Q}}}{2M}\left(1-\frac{2M}{r}\right)\textrm{d}t,\qquad F^{(\alpha,0)}=-\frac{{\mathcal{Q}}}{r^{2}}\,e^{0}\wedge e^{1},\qquad E^{(\alpha,0)}=\frac{{\mathcal{Q}}}{r^{2}}\,e_{1},\qquad B^{(\alpha,0)}=0,$ (2.23)
which corresponds to an electric monopole. While this is the simplest solution one can construct, its electric charge makes it a less appealing model for the magnetosphere. Indeed, the plasma around an electrically charged magnetosphere would quickly discharge any charge excess so that the electrically neutral configuration is the natural assumption when considering the black hole magnetosphere on long timescale. In order to consider a magnetosphere model without electric charge, one can either (i) combine it to the solution $F^{(0,\beta)}$ generated by the azimuthal Killing vector and appropriately choose the combination to make the total electric charge vanishes, this is the well known Wald construction [66], or (ii) one can consider its dual solution $\ast F^{(\alpha,0)}$, which instead has zero electric charge but a non-vanishing magnetic charge.
Let us focus on option (ii), i.e. use the electromagnetic duality of the source-free Maxwell equations and consider the solution given by the Hodge dual of $F^{(\alpha,0)}$:
$\boxed{\tilde{F}^{(\alpha,0)}=\ast F^{(\alpha,0)}=\frac{2a{\mathcal{P}}r\cos{\theta}}{\Sigma^{2}}e^{0}\wedge e^{1}+\frac{{\mathcal{P}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}e^{2}\wedge e^{3}},$ (2.24)
where we have denoted $\mathcal{Q}$} by $\mathcal{P}$}, to stress that this quantity will now represent a magnetic charge. The associated gauge-potential is given by (cf. Notebook 2 in App. A)
$\tilde{A}^{(\alpha,0)}=\tilde{A}^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right).$ (2.25)
Anticipating on the next section, we notice that it satisfies the following interesting property:
$(r^{2}+a^{2})\tilde{A}^{(\alpha,0)}_{t}+a\tilde{A}^{(\alpha,0)}_{\varphi}=0.$ (2.26)
This will reveal useful when studying the general conditions on the electrogeodesic motion. As expected, this dual configuration carries a magnetic charge and no electric charge, i.e. one has
$Q_{\rm e}=\frac{1}{4\pi}\oint_{S}*\tilde{F}^{(\alpha,0)}=0\quad\mbox{and}\quad Q_{\rm m}=\frac{1}{4\pi}\oint_{S}\tilde{F}^{(\alpha,0)}={\mathcal{P}}.$ (2.27)
The electric and magnetic fields measured by the Carter observer are respectively
$\tilde{E}^{(\alpha,0)}$ $=-\frac{2a{\mathcal{P}}r\cos{\theta}}{\Sigma^{2}}\,e_{1}$ (2.28)
$\tilde{B}^{(\alpha,0)}$ $=\frac{{\mathcal{P}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}\,e_{1}.$ (2.29)
They are both purely radial and vanishing for $r\to+\infty$. The solution (2.24) describes thus a magnetic monopole of magnetic charge $\mathcal{P}$} in the Kerr geometry. As we shall see, this solution admits an interesting interpretation in term of the Killing-Yano tensor of the Kerr geometry. In particular, it can be alternatively derived using the Penrose current construction that we discuss in the next section.
Before presenting this construction, let us point that an interesting property of this magnetic monopole solution is that in the non-rotating (Schwarzschild) case, it reduces to
$\tilde{F}^{(\alpha,0)}={\mathcal{P}}\sin{\theta}\,\textrm{d}\theta\wedge\textrm{d}\varphi,$ (2.30)
which is the well known non-rotating Michel solution extensively studied in the literature [67]. This solution describes the simplest magnetic monopole in flat spacetime and on the Schwarzschild black hole. However, in the slow rotating case treated at linear order in $a$, one finds
$\tilde{F}^{(\alpha,0)}\sim{\mathcal{P}}\sin{\theta}\textrm{d}\theta\wedge\left(\textrm{d}\varphi-\frac{a}{r^{2}}\textrm{d}t\right)+\frac{2a{\mathcal{P}}\cos{\theta}}{r^{3}}\textrm{d}t\wedge\textrm{d}r+{\mathcal{O}}(a^{2}),$ (2.31)
which does not correspond to the slowly rotating Michel solution discussed in [67], i.e. $F^{M}=\sin{\theta}\textrm{d}\theta\wedge(\textrm{d}\varphi-\Omega\textrm{d}u)$ where $\textrm{d}u=\textrm{d}t+\textrm{d}r$ and $\Omega$ is a constant. Therefore, the solution (2.24) stands as the non-linear spinning generalization of the non-rotating Michel monopole solution. As we shall see, it enjoys special properties w.r.t to the separability of the electrogeodesic equation.
Asymptotically uniform case: the Wald solution
Let us consider the second branch of solutions which allows one to construct asymptotically uniform magnetic field following Wald’s construction [66]. The second branch corresponds to $\alpha=0$ in Eq. (2.12) and we introduce the constant parameter $\mathcal{B}$} such that
$\beta=\frac{{\mathcal{B}}}{2}.$ (2.32)
There comes then
$A^{(0,\beta)}_{\mu}\textrm{d}x^{\mu}$ $=\beta\chi_{\mu}\textrm{d}x^{\mu}$
$=-\frac{a{\mathcal{B}}Mr\sin^{2}{\theta}}{\Sigma}\,\textrm{d}t+\frac{{\mathcal{B}}\left(r^{4}+a^{2}\Sigma+r^{2}a^{2}\cos^{2}{\theta}+2a^{2}Mr\sin^{2}{\theta}\right)}{2\Sigma}\sin^{2}{\theta}\,\textrm{d}\varphi,$ (2.33)
which yields the following Faraday tensor (cf. Notebook 1 in App. A):
$F^{(0,\beta)}=$ $-\frac{a{\mathcal{B}}\left[M(r^{2}-a^{2}\cos^{2}\theta)+r\Sigma\right]\sin^{2}\theta}{\Sigma^{2}}\,e^{0}\wedge e^{1}-\frac{a{\mathcal{B}}\sqrt{\Delta}\sin\theta}{\Sigma}\,e^{0}\wedge e^{2}+\frac{{\mathcal{B}}\sqrt{\Delta}r\sin\theta}{\Sigma}\,e^{1}\wedge e^{3}$
$+\frac{{\mathcal{B}}\left[r^{4}+a^{2}(r^{2}+2Mr+\Delta\cos^{2}\theta)\right]\cos\theta}{\Sigma^{2}}\,e^{2}\wedge e^{3}.$ (2.34)
Computing the electric and magnetic charges, one finds that this solution carries again only an electric charge, i.e.
$Q_{\rm e}=\frac{1}{4\pi}\oint_{S}\ast F^{(0,\beta)}=2aM{\mathcal{B}}\quad\mbox{and}\quad Q_{\rm m}=\frac{1}{4\pi}\oint_{S}F^{(0,\beta)}=0.$ (2.35)
The electric and magnetic fields measured by the Carter observer are respectively
$E^{(0,\beta)}$ $=\frac{a{\mathcal{B}}\left[M(r^{2}-a^{2}\cos^{2}\theta)+r\Sigma\right]\sin^{2}\theta}{\Sigma^{2}}\,e_{1}+\frac{a{\mathcal{B}}\sqrt{\Delta}\sin\theta}{\Sigma}\,e_{2}$ (2.36)
$B^{(0,\beta)}$ $=\frac{{\mathcal{B}}\left[r^{4}+a^{2}(r^{2}+2Mr+\Delta\cos^{2}\theta)\right]\cos\theta}{\Sigma^{2}}\,e_{1}-\frac{{\mathcal{B}}\sqrt{\Delta}r\sin\theta}{\Sigma}\,e_{2}.$ (2.37)
The first remark is that, at large $r$, while the electric field decays to zero, the magnetic field asymptotes to a value which is independent of $r$, i.e one has
$B^{(0,\beta)}\sim{\mathcal{B}}\left(\cos{\theta}\,e_{1}-\sin{\theta}\,e_{2}\right).$ (2.38)
This provides us with an asymptotically uniform magnetic field of amplitude $\mathcal{B}$} parallel to the rotation axis.
Following the Wald construction [66], it is then natural to consider the linear combination of the two solutions that allows for an electrically neutral solution. Considering the linear combination
$A^{\text{W}}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}+\beta\chi_{\mu}\textrm{d}x^{\mu}$ (2.39)
one finds that the total electric charge is
$Q^{\text{W}}_{e}={\mathcal{Q}}+2aM{\mathcal{B}},$ (2.40)
where $\mathcal{Q}$} and $\mathcal{B}$} are related to $\alpha$ and $\beta$ by Eqs. (2.16) and (2.32). Hence, by choosing ${\mathcal{Q}}=-2aM{\mathcal{B}}$, one obtains an electrically neutral configuration: $Q^{\text{W}}_{e}=0$. This is by far the most studied solution to the Maxwell equations around a Kerr black hole. As we shall see, the electrogeodesic motion is not separable in this case.
Asymptotically growing case: a new toroidal solution
To conclude this section, we present a new way to construct an exact non-vacuum Maxwell solution on the Kerr geometry using the Killing vectors. Consider the following 2-form $F$ that is the Hodge dual of the exterior product of the two Killing 1-forms:
$F_{\mu\nu}=\;\epsilon_{\mu\nu\rho\sigma}\xi^{\rho}\chi^{\sigma}.$ (2.41)
Explicitly, it reads
$F=\;\Sigma\;\textrm{d}r\wedge\textrm{d}\theta\qquad\text{or equivalently}\qquad\boxed{F=\Delta\sin{\theta}\;e^{1}\wedge e^{2}}.$ (2.42)
It is direct to show that $F$ satisfies the non-vacuum Maxwell equations with an electric current given by (cf. Notebook 1 in App. A)
$J^{\mu}\partial_{\mu}=\frac{2}{\Sigma}\left[-\Delta\cos{\theta}\,\partial_{r}+(r-M)\sin{\theta}\,\partial_{\theta}\right].$ (2.43)
In the large $r$ limit, it reduces to $J\sim 2\left(-\cos{\theta}e_{1}+\sin{\theta}e_{2}\right)$, i.e. $J$ becomes uniform and parallel to the rotation axis. With respect to the Carter observer, the electric and magnetic fields read
$E=0\quad\mbox{and}\quad B=\Delta\sin{\theta}\,e_{3}.$ (2.44)
Hence the solution (2.42) corresponds to a purely toroidal magnetic field in the Carter frame. Notice that the field is not uniform but growing at large $r$, so that if it is useful for a magnetosphere model, it can only be considered in a bounded region around the hole. Just as for the Wald solution, the electrogeodesic motion is not separable in this magnetosphere.
Before discussing this point, let us first present the alternative construction of source-free Maxwell solutions based on the so called Penrose current.
Maxwell solutions from Killing-Yano tensors
In the previous section, we have reviewed two approaches to construct vacuum Maxwell solution on the Kerr geometry using the Killing vectors. As it turns out, one can also use the Killing-Yano tensors of the Kerr geometry to build vacuum and non-vacuum Maxwell solutions. As we shall see, some of these solutions can be related to each other using the Killing tower of the Kerr geometry. Therefore, we start by reviewing this key structure ruling the symmetries of the Kerr black hole.
The Killing tower
Consider a metric which admits a tensor satisfying
$\nabla_{\mu}H_{\nu\alpha}=2g_{\mu[\nu}\xi_{\alpha]},\qquad\nabla_{[\mu}H_{\nu\alpha]}=0,\qquad\xi_{\alpha}=\frac{1}{3}\nabla_{\mu}H^{\mu}{}_{\alpha}$ (2.45)
The tensor $H$ is a closed non-generate conformal Killing-Yano (CCKY) 2-form. Following [69, 70, 71], we call this object the principal tensor. Now, taking the dual of the principal tensor, one can build a rank-$2$ antisymmetric tensor satisfing
$Y_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}H^{\rho\sigma},\qquad\nabla_{(\mu}Y_{\nu)\alpha}=0$ (2.46)
which corresponds to a rank-$2$ Killing-Yano (KY) tensor. From this KY tensor, one can construct a rank-$2$ (symmetric) tensor satisfying
$K^{\mu\nu}=Y^{\mu\alpha}Y_{\alpha}{}^{\nu},\qquad\nabla_{(\alpha}K_{\mu\nu)}=0$ (2.47)
which thus provides the Killing tensor for the underlying geometry. Finally, using the principal tensor and the Killing tensor, one can further construct two vectors such that
$\xi^{\mu}\partial_{\mu}$ $=\frac{1}{3}\nabla_{\alpha}H^{\alpha\mu}\partial_{\mu},\qquad\text{such that }\qquad\nabla_{(\mu}\xi_{\nu)}=0$ (2.48)
$\chi^{\mu}\partial_{\mu}$ $=-K^{\mu}{}_{\alpha}\xi^{\alpha}\partial_{\mu},\qquad\text{such that }\qquad\nabla_{(\mu}\chi_{\nu)}=0$ (2.49)
This provides two Killing vectors for the underlying metric. In summary, any metric equipped with a principal tensor is automatically equipped with a rank-$2$ KY tensor and its associated Killing tensor while it also enjoys two isometries. This structure is called the Killing tower, as one can derive all the Killing objects of the metric starting from one single object, the principal tensor. It was identified and described first in the context of higher dimensional black holes in [69, 70, 71]. See [72] for a review.
As it turns out, the Kerr geometry enjoys such a structure. Concretely, the principal tensor and the KY tensor of the Kerr geometry are given by (cf. Notebook 1 in App. A)
$H=re^{0}\wedge e^{1}+a\cos{\theta}\;e^{2}\wedge e^{3}$ (2.50)
$Y=a\cos{\theta}\;e^{0}\wedge e^{1}-re^{2}\wedge e^{3}$ (2.51)
One can easily check that, using the definitions (2.48)–(2.49), one recovers the two commuting Killing vectors
$\xi^{\mu}\partial_{\mu}=\partial_{t}\qquad\chi^{\mu}\partial_{\mu}=\partial_{\varphi}$ (2.52)
Having review these hidden Killing objects of the Kerr geometry, we can now discuss two alternatives ways to construct exact solutions of the Maxwell equations on the Kerr geometry.
The Killing-Maxwell system from the principal tensor
Consider the principal tensor $H$ of the Kerr geometry given by (2.50). It provides an exact solution of the non-vacuum Maxwell equation known as the Killing-Maxwell (KM) solution. To see this, notice first that by definition, it is a closed form so that it satisfies the Bianchi identify $\textrm{d}H=0$. Let us introduce the associated Faraday tensor
$F^{\text{KM}}=\frac{1}{3}H_{\mu\nu}\textrm{d}x^{\mu}\wedge\textrm{d}x^{\nu}=\frac{1}{3}\left(re^{0}\wedge e^{1}+a\cos{\theta}\;e^{2}\wedge e^{3}\right)$ (2.53)
which can be derived from the gauge field
$A^{\text{KM}}=\frac{1}{6}\left[-\left(r^{2}-a^{2}\cos^{2}{\theta}\right)\textrm{d}t+(r^{2}\sin^{2}{\theta}+a^{2}\cos^{2}{\theta})a\textrm{d}\varphi\right]$ (2.54)
Then, using the last equation of (2.45), it can be recast as
$\nabla_{\mu}F_{\text{KM}}^{\mu\nu}=J^{\nu}\qquad\text{with}\qquad J^{\mu}\partial_{\mu}=\partial_{t}$ (2.55)
such that the Maxwell-Gauss equation is satisfied for an electric current given precisely by the time-like Killing vector $\xi^{\mu}\partial_{\mu}=\partial_{t}$. It is remarkable that the fundamental Killing object of the Killing tower provides without any further condition an exact non-vacuum solution of the test Maxwell equation. This system was studied by Carter in [73] where he showed that even more remarkably, it provides one example of a magnetosphere for the Kerr black hole which preserves the existence of the Carter constant for the electrogeodesic motion. The same construction was later studied by Krtous in [61] to build Maxwell solutions on higher dimensional rotating black holes.
The Kerr magnetic monopole from the Penrose current
Another approach to construct vacuum Maxwell solution in a metric possessing a KY tensor was introduced by Penrose in [63]. Consider the following antisymmetric rank-$2$ tensor
$F_{\mu\nu}=\alpha\;R_{\mu\nu\alpha\beta}Y^{\alpha\beta}$ (2.56)
Taking its divergence, one has
$\nabla_{\mu}F^{\mu\nu}$ $=R^{\rho\sigma\mu\nu}\nabla_{\mu}Y_{\rho\sigma}+Y^{\rho\sigma}\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}$ (2.57)
$=-R^{\nu\mu\rho\sigma}\nabla_{[\mu}Y_{\rho\sigma]}+Y^{\rho\sigma}\cancel{\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}}$ (2.58)
$=-R^{\nu[\mu\rho\sigma]}\nabla_{\mu}Y_{\rho\sigma}$ (2.59)
$=0$ (2.60)
Here, we have used the cyclic properties of the Riemann, i.e. $R_{\mu[\nu\rho\sigma]}=0$, the identify $\nabla_{\mu}R_{\rho\sigma}{}^{\mu\nu}=0$ which holds only in vacuum, i.e. when $R_{\alpha\beta}=0$ and the definition of the KY tensor given by
$\nabla_{(\mu}Y_{\nu)\alpha}=0\qquad\nabla_{\mu}Y_{\nu\alpha}=\nabla_{[\mu}Y_{\nu\alpha]}=0\qquad\nabla_{\mu}Y^{\mu}{}_{\alpha}=0$ (2.61)
Taking the Hodge dual, one finally obtains that
$\nabla_{\mu}F^{\mu\nu}=0\qquad\nabla_{\mu}\tilde{F}^{\mu\nu}=0$ (2.62)
where $\tilde{F}_{\mu\nu}=\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ is the dual tensor. Interestingly, applying this to the Kerr geometry and using $\alpha=-{\mathcal{P}}/(2M)$, one obtains [^2] See Notebook 1 in App. A for the computation.
$F$ $=\frac{2a{\mathcal{P}}r\cos{\theta}}{\Sigma^{2}}e^{0}\wedge e^{1}+\frac{{\mathcal{P}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}e^{2}\wedge e^{3},$ (2.63)
which reproduces the solution (2.24) discussed in the previous section. Therefore, the Penrose current of the Kerr spacetime coincides with the dual solution of the Wald construction based solely on the time-like Killing vector $\xi^{\mu}\partial_{\mu}$.
Motion of charged particles in the Kerr magnetosphere
In this section, we review several key properties for the electrogeodesic motion. After presenting the phase space, we discuss the general conditions that the energy and angular momentum of an electrogeodesic have to satisfy in the Kerr geometry. Then, we review the key geometrical property ensuring that the Carter constant is preserved, and thus that the electrogeodesic equations are separable. Finally, we rederive the first order electrogeodesic equations.
In what follows, we shall consider a Maxwell field whose gauge-potential is restricted to the form
$A_{\mu}\textrm{d}x^{\mu}=A_{t}(r,\theta)\,\textrm{d}t+A_{\varphi}(r,\theta)\,\textrm{d}\varphi.$ (3.1)
This covers in particular the case of the test Kerr magnetic monopole which will be the central magnetosphere model we will study.
Lagrangian and phase space
Consider a particle with mass $m$ and electric charge $q$ moving on a curved geometry described by the metric $g$ filled with a gauge-potential $A$. From the beginning, we exclude the case $m=0$ since there are no known charged particles with zero mass. The dynamics of this charged particle is encoded in the Lagrangian
${\mathcal{L}}=\frac{1}{2}mg_{\mu\nu}u^{\mu}u^{\nu}+qA_{\mu}u^{\mu}\qquad\text{with}\qquad u^{\mu}=\frac{\textrm{d}x^{\mu}}{\textrm{d}\lambda}=\dot{x}^{\mu},$ (3.2)
where $\lambda$ is the proper time along the electrogeodesic and we have denoted $x^{\mu}(\lambda)$ the position of the particle and $u^{\mu}=\textrm{d}x^{\mu}/\textrm{d}\lambda$ its $4$-velocity. The equations of motion reads
$mu^{\alpha}\nabla_{\alpha}u^{\mu}=qF^{\mu}{}_{\nu}u^{\nu}.$ (3.3)
The particle is no longer in free fall but moves under the action of both the geometry and the Lorentz force $F^{\mu}{}_{\nu}u^{\nu}$.
Switching to the Hamiltonian formulation, the canonical momenta derived from the Lagrangian is given by
$P_{\mu}=\frac{\delta{\mathcal{L}}}{\delta\dot{x}^{\mu}}=mg_{\mu\nu}\dot{x}^{\nu}+qA_{\mu}.$ (3.4)
It follows that the canonical brackets are given
$\{x^{\mu},x^{\nu}\}=0\qquad\{x^{\mu},P_{\nu}\}=\delta^{\mu}{}_{\nu}\qquad\{P_{\mu},P_{\nu}\}=0$ (3.5)
The Hamiltonian $\mathcal{H}$} takes the form
${\mathcal{H}}=\frac{1}{2}g^{\mu\nu}\left(P_{\mu}-qA_{\mu}\right)\left(P_{\nu}-qA_{\nu}\right)=-\frac{m^{2}}{2}$ (3.6)
From this expression, it is natural to introduce another set of canonical variables and in particular the dressed momenta
$p_{\mu}=P_{\mu}-qA_{\mu}$ (3.7)
which is related to the $4$-velocity of the electrogeodesic flow through $p_{\mu}=mu_{\mu}$. This dressed momenta satisfies the following brackets
$\{x^{\mu},p_{\nu}\}=\delta^{\mu}{}_{\nu}\qquad\{p_{\mu},p_{\nu}\}=qF_{\mu\nu}$ (3.8)
While this momenta is obviously not canonical, it will reveal useful when discussing the conditions to preserve the Carter constant. In term of this momenta, the Hamiltonian takes the usual form
$H=\frac{1}{2}g^{\mu\nu}p_{\mu}p_{\nu}$ (3.9)
such that one has $\{p_{\mu},H\}=0$, i.e. $p_{\mu}$ is indeed conserved along the flow. Therefore, the presence of the Maxwell field introduces a distinction between the canonical momenta $P_{\mu}$, which serves as computing the conserved charges along the flow in GR, and the momenta $p_{\mu}$ associated to the $4$-velocity of the flow $u_{\mu}$. As we shall see, this will have interesting consequences for the trajectories of charged particles. We can now discuss the properties of the electrogeodesics.
General constraints on charged particles
A set of conditions can be derived for the geodesic motion of neutral particles on the Kerr geometry. They descend from the basic requirement that the $4$-velocity $u^{\mu}\partial_{\mu}$ is a time-like vector. They can be expressed in terms of the energy and the angular momentum of the neutral massive particle. As they have important consequences for the geodesic motion, our goal here is to discuss how these constraints are modified for electrogeodesics.
To start with, let us first define the conserved energy $E$ and conserved angular momentum $L$ of the charged particle. As usual, the Kerr geometry being stationary and axisymmetric, one can use its timelike Killing vector $\xi^{\mu}\partial_{\mu}=\partial_{t}$ and the azimuthal Killing vector $\chi^{\mu}\partial_{\mu}=\partial_{\varphi}$ to build $E$ and $L$. Using the canonical momenta (3.4), one obtains
$-E$ $=\xi^{\mu}P_{\mu}=P_{t}=g_{tt}p^{t}+g_{t\varphi}p^{\varphi}+qA_{t}$ (3.10)
$L$ $=\chi^{\mu}P_{\mu}=P_{\varphi}=g_{\varphi\varphi}p^{\varphi}+g_{\varphi t}p^{t}+qA_{\varphi}$ (3.11)
which can be recast into
$E$ $=\left(1-\frac{2Mr}{\Sigma}\right)p^{t}+\frac{2aMr\sin^{2}{\theta}}{\Sigma}p^{\varphi}-qA_{t}$ (3.12)
$L$ $=-\frac{2aMr\sin^{2}{\theta}}{\Sigma}p^{t}+\left(r^{2}+a^{2}+\frac{2a^{2}Mr\sin^{2}{\theta}}{\Sigma}\right)\sin^{2}{\theta}\;p^{\varphi}+qA_{\varphi}$ (3.13)
They provide the corrected notion of energy and momenta in the presence of the Maxwell field.
Let us now review the different conditions on the electrogeodesic.
[…]
    $g(\xi,u)<0,$ (3.14)
    which translates into
    $\frac{P_{t}-qA_{t}}{m}=-\frac{E+qA_{t}}{m}<0.$ (3.15)
    It follows that outside the ergoregion, the energy of the particle is constrained to satisfy
    $\boxed{E>-qA_{t}}.$ (3.16)
    When $q=0$, one recovers the well known result that neutral particles admit only positive energy state outside the ergosphere, i.e. $E>0$ for $r>r_{\rm ergo}(\theta)$.
  • Second constraint : The second constraint can be formulated in term of the scalar product of the vector $N^{\mu}=-\nabla^{\mu}t$ with the particle’s $4$-velocity $u^{\mu}$. Since both vectors are future-directed timelike in the black hole exterior, one must have
    $g(N,u)<0\qquad\rightarrow\qquad u^{t}>0,$ (3.17)
    which can be translated into
    $u^{t}$ $=g^{tt}u_{t}+g^{t\varphi}u_{\varphi}$ (3.18)
    $=\frac{1}{m\Sigma\Delta}\left[\left[(r^{2}+a^{2})\Sigma+2a^{2}Mr\sin^{2}{\theta}\right](E+qA_{t})-2aMr(L-qA_{\varphi})\right]>0.$ (3.19)
    In the region $r_{+}<r<\infty$, one has $\Delta>0$. Therefore, in the outer region of communication of the Kerr geometry, this condition can be simplified to
    $\boxed{(r^{2}+a^{2})\Sigma E-2aMr(L-aE\sin^{2}{\theta})>2qaMr(aA_{t}\sin^{2}{\theta}-A_{\varphi})}.$ (3.20)
  • Third constraint : A third set of conditions can be obtained from the $4$-velocity of the Carter frame, i.e. $e_{0}$. This vector being future-directed timelike its scalar product with $p^{\mu}$ must be negative:
    $g(e_{0},p)<0.$ (3.21)
    Writing $g(e_{0},p)=(e_{0})^{\mu}p_{\mu}$ and using Eq. (2.14a) for $(e_{0})^{\mu}$ and Eq. (3.7) for $p_{\mu}$, with $P_{t}=-E$ and $P_{\varphi}=L$, one gets
    $\boxed{(r^{2}+a^{2})E-aL>-q\left[(r^{2}+a^{2})A_{t}+aA_{\varphi}\right]}.$ (3.22)
    In the case of neutral particles where $q=0$, this condition implies two important consequences. First, for negative energy states with $E\leqslant 0$, which exist only in the ergosphere, the neutral particle must have $L<0$; hence it follows a retrograde orbit. Moreover, for neutral particles with $L=0$, one has automatically that $E>0$. We see that (3.22) does not implies these properties anymore, except if the magnetosphere is such that $(r^{2}+a^{2})A_{t}+aA_{\varphi}=0$. As we shall, see this is the case of the magnetic monopole solution (2.24).
We can now turn to the conditions for separability of the electrogeodesic equations.
Conditions for preserving the Carter constant
An important property of the geodesic motion on the Kerr spacetime is exhibit a set of four constants of motion which are in involution. Three of them are the energy $E$, the angular momentum $L$ which descend from the time and azimuthal isometries of the Kerr geometry. The third one is the mass $m$ of the particle. However, the fourth constant of motion was found by Carter in [54] and descends from the existence of the hidden symmetry of the Kerr geometry giving rise to the Killing tower reviewed in Section 2.2.1. Concretely, for a neutral test particle with momenta $p_{\mu}$, the Killing tensor (2.47) gives rise to a constant ${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu}$ along the geodesic flow, i.e. $D{\mathcal{K}}=0$, known as the Carter constant.
It is therefore natural to wonder under which conditions the Carter constant is preserved for the electrogeodesic motion in a given magnetosphere of the Kerr geometry. This task was explored in [73] and clarified systematically in [74]. See also [75, 76, 77, 78, 79]. Let us recall that the phase space of the electrogeodesic motion can be written as
$\{x^{\mu},p_{\nu}\}=\delta^{\mu}{}_{\nu}\qquad\{p_{\mu},p_{\nu}\}=\kappa F_{\mu\nu}\qquad\text{with}\qquad p_{\mu}=P_{\mu}-qA_{\mu}$ (3.23)
with the Hamiltonian of the form
$H=\frac{1}{2}g^{\mu\nu}p_{\mu}p_{\nu}$ (3.24)
Consider now the following generalized Carter constant defined by
${\mathcal{K}}=K^{\mu\nu}p_{\mu}p_{\nu}$ (3.25)
Then, the Hamiltonian evolution of the function $\mathcal{K}$} is given by
$\{{\mathcal{K}},H\}$ $=\left(\partial_{\rho}K^{\mu\nu}\right)g^{\alpha\beta}p_{\mu}p_{\nu}p_{\beta}\{x^{\rho},p_{\alpha}\}-K^{\mu\nu}\left(\partial_{\rho}g^{\alpha\beta}\right)p_{\nu}p_{\alpha}p_{\beta}\{x^{\rho},p_{\mu}\}+K^{\mu\nu}g^{\alpha\beta}p_{\nu}p_{\beta}\{p_{\mu},p_{\alpha}\}$
$=\frac{1}{6}\nabla^{(\alpha}K^{\mu\nu)}p_{\alpha}p_{\mu}p_{\nu}+F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu}$ (3.26)
which splits into a term cubic in the momenta and a second term quadratic in the momenta. By definition of the Killing tensor (2.47), the first term vanishes reducing the bracket to
$\{{\mathcal{K}},H\}$ $=F_{\mu}{}^{(\alpha}K^{\nu)\mu}p_{\alpha}p_{\nu}$ (3.27)
which can be interpreted as the anomaly induced by the presence of the magnetosphere. Therefore, the condition for the Carter constant to remain a constant of motion of the electrogeodesic motion is given by
$\{{\mathcal{K}},H\}=0\qquad\text{if}\qquad F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0$ (3.28)
Since the Killing tensor descends from the KY $2$-form of the Kerr geometry, this condition can be written as follows
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\alpha}Y^{\beta\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\nu}Y^{\beta\mu})Y^{\alpha}_{\beta}$ (3.29)
Let us now impose that
$S^{\alpha\beta}=F_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0$ (3.30)
Plugging this relation in the previous expression, one finds
$F_{\mu}{}^{(\alpha}K^{\nu)\mu}$ $=(F_{\mu}{}^{\beta}Y^{\alpha\mu})Y^{\nu}{}_{\beta}+(F_{\mu}{}^{\beta}Y^{\nu\mu})Y^{\alpha}_{\beta}=Y^{\alpha\beta}Y^{\nu\mu}\left(F_{\beta\mu}+F_{\mu\beta}\right)=0$ (3.31)
Therefore, the condition (3.30) is sufficient to ensures that $F_{\mu}{}^{(\alpha}K^{\nu)\mu}=0$, as first noticed in [74]. This condition selects the models of Kerr magnetosphere which preserve the Carter constant.
We can now check which solution for the Kerr magnetosphere satisfies this peculiar condition. Let us examine each magnetosphere models discussed in the previous section.
  • Killing-Maxwell solution: It is direct to check that
    $F_{\alpha\beta}=H_{\alpha\beta},\qquad S^{\alpha\beta}=H_{\mu}{}^{[\alpha}Y^{\beta]\mu}=0$ (3.32)
    To our knowledge, together with the Kerr-Newman black hole magnetosphere, this is the only Maxwell solution satisfying that property. This was shown by Carter in [73].
  • The Kerr magnetic monopole (and Penrose current): Consider now the magnetosphere built from the Penrose current (or the dual time-like Killing vector solution). One has
    $F_{\mu\nu}=R_{\mu\nu\rho\sigma}Y^{\rho\sigma}\qquad S^{\alpha\beta}=R_{\mu}{}^{[\alpha}{}_{\rho\sigma}Y^{\rho\sigma}Y^{\beta]\mu}=0$ (3.33)
    This is expected as this magnetosphere is the magnetic subsector of the Kerr-Newman magnetosphere. This opens the door to analytically study the electrogeodesic motion in this Kerr magnetosphere which is considered test, i.e. without backreaction. This will be discussed in the next section.
  • Wald solution: Consider now the Wald solution. One finds [^4] See Notebook 1 in App. A for the computation. that $S^{\alpha\beta}\neq 0$ in this case: the 2-form $\underline{S}=S_{\alpha\beta}\;\textrm{d}x^{\alpha}\wedge\textrm{d}x^{\beta}$ metric-dual to $S^{\alpha\beta}$ is
    $\underline{S}=\frac{{\mathcal{B}}}{2}\sqrt{\Delta}\sin\theta\,e^{1}\wedge e^{2}.$ (3.34)
Having reviewed the conditions for a magnetosphere to preserve the Carter constant, and demonstrated that the Kerr monopole solution belongs to this very special class, we naturally turn towards writing the first order electrogeodesic equations.
Electrogeodesic motion in the Kerr monopole magnetosphere: Separability
The main goal of this section is to review the different steps to separate of the electrogeodesic equations within the Kerr test monopole magnetosphere. The first order electrogeodesic equations we shall derive are a subset of the electrogeodesic equations of the Kerr-Newman black hole [55] where one set the electric charge to zero and keep the magnetic charge arbitrary.
Separability of the radial and the polar motion
In order to separate the radial equation, we first consider first the Carter constant given by
${\mathcal{K}}=K_{\mu\nu}p^{\mu}p^{\nu}$ (3.35)
where $K_{\mu\nu}$ is the Killing tensor of the Kerr geometry which can be decompose as
$K_{\alpha\beta}=(r^{2}+a^{2})(k_{\alpha}\ell_{\beta}+\ell_{\alpha}k_{\beta})+r^{2}g_{\alpha\beta}$ (3.36)
Here $k^{\alpha}\partial_{\alpha}$ and $\ell^{\alpha}\partial_{\alpha}$ are the null vectors corresponding to the principal null directions of Kerr geometry. They are given explicitly by
$k^{\alpha}\partial_{\alpha}$ $=\frac{r^{2}+a^{2}}{\Delta}\partial_{t}-\partial_{r}+\frac{a}{\Delta}\partial_{\varphi}$ (3.37)
$\ell^{\alpha}\partial_{\alpha}$ $=\frac{1}{2}\partial_{t}+\frac{\Delta}{2(r^{2}+a^{2})}\partial_{r}+\frac{a}{2(r^{2}+a^{2})}\partial_{\varphi}$ (3.38)
A direct computation of the Carter constant gives
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})(E+qA_{t})-a(L-qA_{\varphi})\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.39)
$=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL+{q((r^{2}+a^{2})A_{t}+aA_{\varphi})}\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.40)
such that one obtains a non-trivial contribution from the magnetosphere encoded in the last term. Interestingly, focusing on the Kerr monopole, one finds that this term vanishes,
$(r^{2}+a^{2})A_{t}+aA_{\varphi}=0$ (3.41)
so that the Carter constant takes the same form as for geodesic
${\mathcal{K}}$ $=\frac{1}{\Delta}\left[\left((r^{2}+a^{2})E-aL\right)^{2}-\Sigma^{2}(p^{r})^{2}\right]-r^{2}m^{2}$ (3.42)
Since this constant depends only on the radial component $p^{r}$, it allows one to algebraically deduce the first order radial equation which reads
$\Sigma\;p^{r}=\Sigma m\frac{\textrm{d}r}{\textrm{d}\lambda}=\sqrt{R(r)}\qquad\text{with}\qquad R(r)=\left[(r^{2}+a^{2})E-aL\right]^{2}-\Delta(r^{2}m^{2}+{\mathcal{K}})$ (3.43)
This equation reproduces the one of the geodesic case, but one has to keep in mind that now, it involves a corrected notion of the energy and angular momentum. Notably, the l.h.s only involves functions of the radius $r$.
Now, we can obtain the first order equation for the polar motion using both the Carter constant and the mass relation $m^{2}=-g_{\mu\nu}p^{\mu}p^{\nu}$. Writing this explicitly, one finds that
$m^{2}=-g^{\mu\nu}\left(P_{\mu}-qA_{\mu}\right)\left(P_{\nu}-qA_{\mu}\right)=-g^{\mu\nu}P_{\mu}P_{\nu}+2qg^{\mu\nu}P_{\mu}A_{\nu}-q^{2}g^{\mu\nu}A_{\mu}A_{\nu}$ (3.44)
Since $A_{r}=A_{\theta}=0$ so that $P_{r}=p_{r}$ and $P_{\theta}=p_{\theta}$, the first term will reproduce the standard term for the geodesic. The challenge lies in the two last terms which need to provide separable form. Concretely, they are given by
$g^{\mu\nu}P_{\mu}A_{\nu}$ $=-E\left(g^{tt}A_{t}+g^{t\varphi}A_{\varphi}\right)+L\left(g^{t\varphi}A_{t}+g^{\varphi\varphi}A_{\varphi}\right)=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(aE-\frac{L}{\sin^{2}{\theta}}\right)$ (3.45)
$g^{\mu\nu}A_{\mu}A_{\nu}$ $=\frac{{\mathcal{P}}^{2}}{\Sigma}\frac{\cos^{2}{\theta}}{\sin^{2}{\theta}}$ (3.46)
Equation (3.44) becomes then
$m^{2}$ $=-\frac{1}{\Delta}\left(r^{2}+a^{2}+\frac{2a^{2}Mr\sin^{2}{\theta}}{\Sigma}\right)E^{2}-\frac{4aMr}{\Sigma\Delta}EL-\frac{1}{\Delta}\left(1-\frac{2Mr}{\Sigma}\right)\frac{L^{2}}{\sin^{2}{\theta}}$
$\qquad+\frac{2q{\mathcal{P}}\cos{\theta}}{\Sigma}\left(aE-\frac{L}{\sin^{2}{\theta}}\right)-\frac{q^{2}{\mathcal{P}}^{2}}{\Sigma}\frac{\cos^{2}{\theta}}{\sin^{2}{\theta}}$
$\qquad-\frac{\Sigma}{\Delta}(p^{r})^{2}-\Sigma(p^{\theta})^{2}$ (3.47)
Upon using (3.43) to rewrite $p^{r}$ as
$\frac{\Sigma}{\Delta}(p^{r})^{2}=\frac{1}{\Delta}\left[(r^{2}+a^{2})E-aL\right]^{2}-(r^{2}m^{2}+{\mathcal{K}})$ (3.48)
and injecting in Eq. (3.47), one obtains after a little algebra that
$\Sigma\;p^{\theta}=\Sigma m\frac{\textrm{d}\theta}{\textrm{d}\lambda}=\pm\sqrt{\Theta(\theta)}$ (3.49)
with the potential given by
$\Theta(\theta)={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos\theta}{\sin{\theta}}-aE\sin{\theta}\right)^{2}$ (3.50)
The polar motion follows the same equation as for the geodesic which also involves only functions of the polar angle $\theta$. In conclusion, the property (3.30) provides us with a Carter-like constant which, together with the mass, can be used to derive the first order equations for the radial and polar motion. Both equations are separable and remain unchanged w.r.t to the geodesic motion. This already implies that the different radial and polar equilibrium positions for the charged particles will remain the same as the one of their neutral counterpart. We shall come back on this point in the following.
Separability of the equation for the time and azimuthal motion
Let us start with the equations for $\dot{t}$ and $\dot{\varphi}$. The expressions for the energy and the angular momentum can be written as
$\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right)$ $=\left(\begin{array}[]{cccc}g_{tt}&g_{t\varphi}\\ g_{\varphi t}&g_{\varphi\varphi}\end{array}\right)\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right)$ (3.57)
Inverting this system, one obtains
$\left(\begin{array}[]{cc}p^{t}\\ p^{\varphi}\end{array}\right)=\frac{1}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}\left(\begin{array}[]{cccc}g_{\varphi\varphi}&-g_{t\varphi}\\ -g_{\varphi t}&g_{tt}\end{array}\right)\left(\begin{array}[]{cc}-E-qA_{t}\\ L-qA_{\varphi}\end{array}\right)$ (3.58)
which provides us with the two following equations
$m\frac{\textrm{d}t}{\textrm{d}\lambda}$ $=-\frac{g_{t\varphi}L+g_{\varphi\varphi}E}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}-q\frac{g_{\varphi\varphi}A_{t}-g_{t\varphi}A_{\varphi}}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}$ (3.59)
$m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{g_{tt}L+g_{t\varphi}E}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}-q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}$ (3.60)
When the magnetosphere is turned off, i.e. when $q=0$, one recovers the standard equations for the time and azimuthal motions of the neutral particles on the Kerr geometry. They are well known to be separable in term of the Mino time $\textrm{d}\tau=\textrm{d}\lambda/\Sigma$. The key challenge when $q\neq 0$ is that the new terms do not spoil this separability. Concretely, upon factorizing by the function $1/\Sigma$, they should appear as functions of the radius $r$ or of the polar angle $\theta$ only. Notice that this is not guaranteed by the presence of the Carter constant. However, it turns out to be the case. To see this, one can use that
$g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}=-\Delta\sin^{2}{\theta}$ (3.61)
to show that
$q\frac{g_{\varphi\varphi}A_{\varphi}-g_{t\varphi}A_{t}}{-\Delta\sin^{2}{\theta}}$ $=-q\frac{a{\mathcal{P}}\cos{\theta}}{\Sigma}$ (3.62)
$q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{-\Delta\sin^{2}{\theta}}$ $=-q\frac{{\mathcal{P}}\cos{\theta}}{\Sigma\sin^{2}{\theta}}$ (3.63)
Upon an overall factor of $1/\Sigma$, both contributions only involve a function of the polar angle. This ensures that the equations for $\dot{t}$ and $\dot{\varphi}$ are indeed separable. They are explicitly given by
$\Sigma m\frac{\textrm{d}t}{\textrm{d}\lambda}$ $=\frac{1}{\Delta}\left[(r^{2}+a^{2})^{2}E-2MarL\right]-a^{2}E\sin^{2}{\theta}+aq{\mathcal{P}}\cos{\theta}$ (3.64)
$\Sigma m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{L}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right)+\frac{q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}$ (3.65)
This concludes the review of the proof of the separability of the electrogeodesic equations in the test magnetic monopole on the Kerr geometry.
Integral solution to the electrogeodesic equation
We are now in position to provide the exact analytic solution to the electrogeodesic equation. For clarity, let us gather the four equations. Introducing the Mino time
$\textrm{d}\tau=\frac{\textrm{d}\lambda}{\Sigma(r(\lambda),\theta(\lambda))}$ (3.66)
the system is given by
$m\frac{\textrm{d}t}{\textrm{d}\tau}$ $=\frac{1}{\Delta}\left[(r^{2}+a^{2})^{2}E-2aMrL\right]-a^{2}E\sin^{2}{\theta}+aq{\mathcal{P}}\cos{\theta}$ (3.67a)
$m\frac{\textrm{d}r}{\textrm{d}\tau}$ $=\epsilon_{r}\sqrt{R(r)}$ (3.67b)
$m\frac{\textrm{d}\theta}{\textrm{d}\tau}$ $=\epsilon_{\theta}\sqrt{\Theta(\theta)}$ (3.67c)
$m\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right)$ (3.67d)
where $\epsilon_{r}=\pm 1$, $\epsilon_{\theta}=\pm 1$ and the radial and polar potentials are given by
$R(r)$ $=\left[(r^{2}+a^{2})E-aL\right]^{2}-\Delta(r^{2}m^{2}+{\mathcal{K}})$ (3.68)
$\Theta(\theta)$ $={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-aE\sin{\theta}\right)^{2}$ (3.69)
Let us introduce the specific conserved energy $\varepsilon$, the specific conserved angular momentum $\ell$, the reduced Carter constant $k$ and the charge-to-mass ratio $\kappa$ by
$\varepsilon=\frac{E}{m},\quad\ell=\frac{L}{m},\quad k=\frac{{\mathcal{K}}}{m^{2}}\quad\mbox{and}\quad\kappa=\frac{q}{m}.$ (3.70)
Note that, in the geometrical ($G=1$ and $c=1$) and Gaussian units that we are using, $\varepsilon$ and $\kappa$ are dimensionless, $\ell$ has the dimension of a length and $k$ that of a squared length. In terms of the reduced quantities (3.70), the system (3.67) becomes
$\frac{\textrm{d}t}{\textrm{d}\tau}$ $=\frac{1}{\Delta}\left[(r^{2}+a^{2})^{2}\varepsilon-2aMr\ell\right]-a^{2}\varepsilon\sin^{2}{\theta}+a\kappa{\mathcal{P}}\cos{\theta}$ (3.71a)
$\frac{\textrm{d}r}{\textrm{d}\tau}$ $=\epsilon_{r}\sqrt{{\mathcal{R}}(r)}$ (3.71b)
$\frac{\textrm{d}\theta}{\textrm{d}\tau}$ $=\epsilon_{\theta}\sqrt{\tilde{\Theta}(\theta)}$ (3.71c)
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2Mr\varepsilon-a\ell\right)$ (3.71d)
with
${\mathcal{R}}(r)$ $=\left[(r^{2}+a^{2})\varepsilon-a\ell\right]^{2}-\Delta(r^{2}+k)$ (3.72)
$\tilde{\Theta}(\theta)$ $=k-a^{2}\cos^{2}{\theta}-\left(\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-a\varepsilon\sin{\theta}\right)^{2}$ (3.73)
Note that ${\mathcal{R}}(r)=R(r)/m^{2}$ and $\tilde{\Theta}(\theta)=\Theta(\theta)/m^{2}$ and the scaling have been chosen so that the particle’s mass $m$ does not appear in the above formulas. Note as well that $\kappa{\mathcal{P}}$ has the dimension of a length.
The radial and polar equations (3.71b) and (3.71c) imply that
$\tau-\tau_{0}$ $=\int^{r}_{r_{0}}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}=\int^{\theta}_{\theta_{0}}\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}$ (3.74)
Then, the time and azimuthal equations (3.71a) and (3.71d) can be integrated as
$t-t_{0}$ $=\int^{r}_{r_{0}}\frac{(\bar{r}^{2}+a^{2})^{2}\varepsilon-2aM\bar{r}\ell}{\Delta(\bar{r})}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}-\int^{\theta}_{\theta_{0}}\left[a^{2}\varepsilon\sin^{2}{\bar{\theta}}-a\kappa{\mathcal{P}}\cos{\bar{\theta}}\right]\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}$ (3.75)
$\varphi-\varphi_{0}$ $=a\int^{r}_{r_{0}}\frac{2M\bar{r}\varepsilon-a\ell}{\Delta(\bar{r})}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}+\int^{\theta}_{\theta_{0}}\left[\frac{\ell+\kappa{\mathcal{P}}\cos{\bar{\theta}}}{\sin^{2}{\bar{\theta}}}\right]\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}$ (3.76)
These integrals formulas provide a closed form solution for the electrogeodesic motion in the Kerr monopole magnetopshere. As expected, the correction terms proportional to $\kappa$ being purely polar dependent, they can be integrated out without further difficulty. We can now use this result to study the ejection of charged particles in this specific magnetosphere.
Jet launching in the vacuum Kerr monopole
In this section, we use the exact solution to the electrogeodesic equation derived in the previous section to study the acceleration of charged particles near the poles and the formation of jets. We present the exact formula for the acceleration. Then, we classify the stable and unstable polar equilibrium positions which define the opening angle of the jet at its basis. We further derive the expression for the azimuthal speed experienced by the charged particles in this jet. Finally, we compute the electromagnetic gravitational redshift measured by an asymptotic Carter observer and determine the condition under which the ejected charged particle is blueshifted. To our knowledge, the only other investigations of the motion of charged particles in the Kerr monopole have been discussed in [55] and more recently in [50].
We have seen that the electrogeodesics satisfies general conditions given by (3.16), (3.20) and (3.22). Applied to the magnetic monopole magnetosphere given by [cf. Eq. (2.25)]
$A_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right)$ (4.1)
conditions (3.16) and (3.22) respectively give
$\varepsilon>-\frac{a\kappa{\mathcal{P}}\cos{\theta}}{\Sigma}\quad\text{for}\quad r>r_{\rm ergo}(\theta)$ (4.2)
$\varepsilon>\Omega_{\rm C}\ell,$ (4.3)
where we have used property (3.41) and have let appear the angular velocity $\Omega_{\rm C}=\Omega_{\rm C}(r)$ of the Carter observer, via $\Omega_{\rm C}=a/(r^{2}+a^{2})$ [Eq. (2.13)]. Along the Northern rotation axis ($\theta=0$), where $r_{\rm ergo}(0)=r_{+}$, we may rewrite (4.2) as
$\varepsilon>-\Omega_{\rm C}\kappa{\mathcal{P}}\qquad(\theta=0).$ (4.4)
The above conditions will serve as consistency checks for the following discussion.
Acceleration of charged particles
With this solution at end, one can obtain an analytic formula for the 4-acceleration $a^{\mu}$ of the charged particles, which reads
$a^{\mu}=\kappa F^{\mu}{}_{\nu}u^{\nu}.$ (4.5)
It is easier to perform the computation in the Carter tetrad $(e_{\alpha})$, as given by Eq. (2.14). The components of the particle’s 4-velocity with respect to the coordinate frame are $u^{\alpha}=\Sigma^{-1}\textrm{d}x^{\alpha}/\textrm{d}\tau$ with $\textrm{d}x^{\alpha}/\textrm{d}\tau$ given by the electrogeodesic system (3.71). Transforming to the Carter tetrad via the system (2.14), we get
$u=\frac{\varepsilon(r^{2}+a^{2})-a\ell}{\sqrt{\Sigma}\sqrt{\Delta}}\,e_{0}+\epsilon_{r}\frac{\sqrt{{\mathcal{R}}(r)}}{\sqrt{\Sigma}\sqrt{\Delta}}\,e_{1}+\epsilon_{\theta}\frac{\sqrt{\tilde{\Theta}(\theta)}}{\sqrt{\Sigma}}\,e_{2}+\frac{\ell+\kappa{\mathcal{P}}\cos\theta-a\varepsilon\sin^{2}\theta}{\sqrt{\Sigma}\sin\theta}\,e_{3}.$ (4.6)
Using the components (2.24) of the Faraday tensor in the Carter tetrad, Eq. (4.5) leads to the particle’s 4-acceleration in the Carter frame [^5] See Notebook 2 in App. A for the computation.:
$a^{0}=$ $-\epsilon_{r}\frac{2a\kappa{\mathcal{P}}r\cos\theta\sqrt{{\mathcal{R}}(r)}}{\Sigma^{5/2}\sqrt{\Delta}}$ (4.7a)
$a^{1}=$ $-\frac{2a\kappa{\mathcal{P}}\left[(r^{2}+a^{2})\varepsilon-a\ell\right]r\cos\theta}{\Sigma^{5/2}\sqrt{\Delta}}$ (4.7b)
$a^{2}=$ $\frac{\kappa{\mathcal{P}}(\ell+\kappa{\mathcal{P}}\cos\theta-a\varepsilon\sin^{2}\theta)(r^{2}-a^{2}\cos^{2}\theta)}{\Sigma^{5/2}\sin\theta}$ (4.7c)
$a^{3}=$ $-\epsilon_{\theta}\frac{\kappa{\mathcal{P}}\sqrt{\tilde{\Theta}(\theta)}(r^{2}-a^{2}\cos^{2}\theta)}{\Sigma^{5/2}}.$ (4.7d)
The norm of the 4-acceleration is
$\sqrt{a_{\mu}a^{\mu}}=\frac{|\kappa{\mathcal{P}}|}{\Sigma^{3/2}}\sqrt{k+\frac{a^{2}}{\Sigma}(3r^{2}-a^{2}\cos^{2}\theta)\cos^{2}\theta}.$ (4.8)
Note that it involves only $|\kappa{\mathcal{P}}|$ and the reduced Carter constant $k$. Let us now investigate what these analytic expressions teach us about the 4-acceleration experienced by the charged particles in the different regions.
Consider first the equatorial plane, i.e. $\theta=\pi/2$. The components $a^{0}$ and $a^{1}$ of the 4-acceleration are proportional to $\cos{\theta}$, such that they vanish identically at the equator:
$a^{0}(r,\pi/2)=a^{1}(r,\pi/2)=0$ (4.9)
As for the azimuthal and polar components, one finds that
$a^{2}(r,\pi/2)$ $=\frac{\kappa{\mathcal{P}}(\ell-a\varepsilon)}{r^{1/2}}$ (4.10)
$a^{3}(r,\pi/2)$ $=-\epsilon_{\theta}\kappa{\mathcal{P}}\frac{\sqrt{k-(\ell-a\varepsilon)^{2}}}{r^{1/2}}$ (4.11)
We see that for a general electrogeodesic labelled by $(\varepsilon,\ell,k)$, the polar acceleration will drive it away from the equatorial plane, suggesting that this position is no longer an equilibrium position as for the geodesic case. Yet, an exception remains: it corresponds to the electrogeodesics for which $\ell=a\varepsilon$ (note that this satisfies the inequality (4.3)). In this case, the polar acceleration vanishes, i.e. $a^{2}(r,\pi/2)=0$, potentially allowing for an equilibrium position. This can be confirmed by the analysis of the polar potential which is the subject of the next subsection.
On the contrary, near the poles, at $\theta=0$ (or equivalently $\theta=\pi$), the time and radial components of the acceleration $(a^{0},a^{1})$ take their maximal value, which suggests that charged particles are ejected preferentially near the poles. Let us also point that thanks to condition (4.3), the term $(r^{2}+a^{2})\varepsilon-a\ell$ in the radial acceleration (4.7b) cannot vanish.
There is a location at which the particle’s 4-acceleration $a^{\mu}$ vanishes. Since $a^{\mu}$ is necessarily a spacelike vector, its vanishing is equivalent to its norm $\sqrt{a_{\mu}a^{\mu}}$ being zero. From expression (4.8) of the latter, we get the condition
$r=r_{\ast}(\theta)=a|\cos{\theta}|\sqrt{\frac{a^{2}\cos^{2}\theta-k}{3a^{2}\cos^{2}\theta+k}}.$ (4.12)
Recall that $k\geq 0$; hence we must have a not too large Carter constant: $0\leq k\leq a^{2}\cos^{2}\theta$. At this radius, the electrogeodesics reduces to a free-fall motion. Properly characterizing the role of this boundary in the magnetosphere will require more work which will be presented elsewhere. However, we already anticipate that this locus could play a role in allowing spherical non-accelerated trajectories.
Finally, it is interesting to obtain the behavior of the acceleration at large $r$ which reads
$a^{0}\sim a^{(t)}\sim$ $\;-2\epsilon_{r}a\kappa{\mathcal{P}}\sqrt{\varepsilon^{2}-1}\,\frac{\cos\theta}{r^{3}}$ (4.13a)
$a^{1}\sim a^{(r)}\sim$ $\;-2a\kappa{\mathcal{P}}\varepsilon\,\frac{\cos\theta}{r^{3}}$ (4.13b)
$a^{2}\sim a^{(\theta)}\sim$ $\;\frac{\kappa{\mathcal{P}}(\ell+\kappa{\mathcal{P}}\cos\theta-a\varepsilon\sin^{2}\theta)}{r^{3}\sin\theta}$ (4.13c)
$a^{3}\sim a^{(\varphi)}\sim$ $\;-\epsilon_{\theta}\frac{\kappa{\mathcal{P}}\sqrt{\tilde{\Theta}(\theta)}}{r^{3}},$ (4.13d)
while the the norm behaves as
$\sqrt{a_{\mu}a^{\mu}}\sim\frac{|\kappa{\mathcal{P}}|}{r^{3}}\sqrt{k+3a^{2}\cos^{2}\theta}.$ (4.14)
This concludes our discussion on the acceleration. It is crucial to notice that the present analysis has been made possible thanks to the exact solution of the electrogeodesic motion. As we can see, the expression for the acceleration experienced by the test charged particles already reveals several interesting features of the magnetosphere and the motion it induces on the particles. The key point is that even in this vacuum magnetosphere, the charged particles are naturally driven towards the poles and preferentially ejected near the poles.
Structure of the jet
We can now study how to characterize the jet. To this end, it is useful to identify the equilibrium positions of the polar motion and the stable versus unstable nature. Since the idealized jet corresponds to the charged particles escaping at infinity along the rotation axis, we first focus on characterizing the conditions for the pole to be a stable position. As we shall see, the presence of the magnetosphere introduces again two regimes depending on the energy of the charged particles. Then, we further discuss the case of the equatorial plane.
Stable versus unstable polar equilibrium positons
Consider the charged particles which can cross the rotation axis located $\theta=0$ in the Northern hemisphere. From expression (3.73) of $\tilde{\Theta}(\theta)$, this requires
$\ell=-\kappa{\mathcal{P}}.$ (4.15)
to cancel the divergence in the potential. Notice that in the neutral case, one recovers the standard result that $\ell=0$. Moreover, for the value (4.15) of $\ell$ the general condition (4.3), i.e. $\varepsilon>\Omega_{\rm C}\ell$, reduces to condition (4.4), namely $\varepsilon>-\Omega_{\rm C}\kappa{\mathcal{P}}$.
Sticking to the value (4.15) of $\ell$ and assuming $a\neq 0$, the latitudinal equation of motion (3.71c) can be rewritten as
$\left(\frac{\textrm{d}\theta}{\textrm{d}\tau}\right)^{2}+V(\theta)=k$ (4.16)
with the effective potential [^6] Use has been made of the identity $(1-\cos\theta)/\sin\theta=\tan(\theta/2)$.
$V(\theta)$ $=a^{2}\left[\cos^{2}\theta+\left(\varepsilon\sin\theta-\beta\tan\frac{\theta}{2}\right)^{2}\right]$ (4.17)
where $\beta$ is the dimensionless quantity defined by
$\beta=-\frac{\kappa{\mathcal{P}}}{a}=\frac{\ell}{a}.$ (4.18)
As we shall see, this parameter is crucial to characterize the new features induced by the magnetosphere. In term of $\beta$, condition (4.3) reads
$\varepsilon>a\Omega_{C}\beta.$ (4.19)
By definition, the equilibrium polar positions for the charged particles satisfy
$V(\theta)=0\qquad V^{\prime}(\theta)=0$ (4.20)
and it is stable by $V^{\prime\prime}(\theta)>0$ and unstable if $V^{\prime\prime}(\theta)<0$. The effective potential $V(\theta)$ is depicted in Figs. 12.
Let us first consider the north pole $\theta=0$ which corresponds to the idealized jet, i.e. to charged particles escaping to infinity along the rotation axis. In this case, one finds
$V(0)$ $=a^{2},\qquad V^{\prime}(0)=0,\qquad V^{\prime\prime}(0)=2a^{2}(\varepsilon-\beta/2-1)(\varepsilon-\beta/2+1)$ (4.21)
It follows that $\theta=0$ is a stable constant latitude position, i.e. that $\textrm{d}\theta/\textrm{d}\tau=0$ and $V^{\prime\prime}(0)>0$, if and only if, the electrogeodesic $(\varepsilon,\ell,k)$ satisfies
$k$ $=a^{2}\qquad\text{and}\qquad\varepsilon<\frac{\beta}{2}-1\quad\mbox{or}\quad\varepsilon>1+\frac{\beta}{2}.$ (4.22)
Taking into account condition (4.19), one obtains the following conditions:
$a\Omega_{\rm C}\beta<\varepsilon<\frac{\beta}{2}-1\qquad\mbox{or}\qquad\varepsilon>\frac{\beta}{2}+1.$ (4.23)
For the second condition, we have taken into account that $a\Omega_{\rm C}=a^{2}/(r^{2}+a^{2})$ always fulfils $a\Omega_{\rm C}<1/2$, with the upper bound $1/2$ approached only near the event horizon of an extremal ($a=M$) Kerr black hole, so that $\varepsilon>a\Omega_{C}\beta$ is actually implied by $\varepsilon>{\beta}/{2}+1$.
It follows that there are two distinct families of electrogeodesics defined by the above thresholds which corresponds to the idealized jet. The expression of $V^{\prime}(\theta)$ is
$V^{\prime}(\theta)=\frac{2a^{2}}{\sin^{2}\theta}\left\{2\beta^{2}\tan\frac{\theta}{2}+\sin\theta\left[\left((\varepsilon^{2}-1)\cos\theta-\varepsilon\beta\right)\sin^{2}\theta-\beta^{2}\right]\right\}$ (4.24)
but note that the potential $V(\theta)$ is meaningful only for $\ell=-\kappa{\mathcal{P}}$.
Refer to caption Refer to caption
Figure 1: Effective potential $V(\theta)$ governing the latitudal motion for $\ell=-\kappa{\mathcal{P}}$. The plots are performed for $\beta=8$ and for various values of $\varepsilon$ ensuring a stable equilibrium at $\theta=0$, according to Eq. (4.22), i.e. $\varepsilon<3$ (left panel) or $\varepsilon>5$ (right panel). Notice that in all cases, $\theta=0$ is a (local) minimum of $V$ and that the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots.
Figure 2: Effective potential $V(\theta)$ governing the latitudal motion for $\ell=-\kappa{\mathcal{P}}$ and $\beta=8$. Contrary to Fig. 1, the chosen values of $\varepsilon$ are such that the equilibrium position at $\theta=0$ is unstable: it corresponds to a local maximum of $V(\theta)$. According to Eq. (4.22) with $\beta=8$, these values are in the range $3<\varepsilon<5$ (the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots). Notice that a stable equilibrium position occurs at $\theta=\theta_{*}\in(0,\pi/2)$.
To conclude this section, let us further comment on the parameter $\beta$. First, notice that for an electron [^7] In terms of SI units, one has actually $\kappa=\sqrt{\mu_{0}/(4\pi)}\,q/m$, so that, for an electron, $\kappa_{\rm e}=-5.56\;10^{7}\;{\rm m}^{1/2}{\rm kg}^{-1/2}$, which yields $\kappa_{\rm e}=-2.04\;10^{21}$ in geometrical units., $\kappa=\kappa_{\rm e}=-2.04\;10^{21}$, while for a proton, $\kappa=\kappa_{\rm p}=1.11\;10^{18}$. These huge values reflect the large dominance of the electromagnetic interaction over the gravitational one for elementary particles. On the contrary, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$, in order for the electromagnetic field to not significantly modify the spacetime metric, which is assumed to be the Kerr metric; otherwise one would have to consider the Kerr-Newman metric. For values of $a\sim M$, this is equivalent to demand $|{\mathcal{P}}|/a\ll 1$. It follows that $\beta$ can take values of order one or even much higher. For instance, for electrons, for which $|\kappa|\sim 10^{21}$, $\beta$ can be as large as $10^{10}$ for $|{\mathcal{P}}|=10^{-10}M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
Azimuthal speed and magnetic frame-dragging
Let us now discuss the azimuthal angular velocity experienced by the charged particles. Imposing $\ell=-\kappa{\mathcal{P}}$, one obtains
$\frac{\textrm{d}t}{\textrm{d}\tau}$ $=\frac{1}{\Delta}\left[(r^{2}+a^{2})^{2}\varepsilon+2\kappa aM{\mathcal{P}}r\right]-a^{2}\varepsilon\sin^{2}{\theta}+a\kappa{\mathcal{P}}\cos{\theta}$ (4.25)
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{a}{\Delta}\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)+\frac{\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)}{\sin^{2}{\theta}}$ (4.26)
The angular velocity is then given by
$\Omega(r,\theta)=\frac{\textrm{d}\varphi}{\textrm{d}t}=\frac{\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)a\sin^{2}{\theta}+\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)\Delta}{\left[(r^{2}+a^{2})^{2}\varepsilon+2\kappa aM{\mathcal{P}}r+a\Delta\left(\kappa{\mathcal{P}}\cos{\theta}-a\varepsilon\sin^{2}{\theta}\right)\right]\sin^{2}{\theta}}$ (4.27)
As expected, this angular velocity is not defined for particles located at the poles, i.e. at $\theta=0$. However, this formula provides the angular velocity of the particle for $\theta\in\;]0,\pi/2]$.
Let us now evaluate its expression in the large $r$ limit. Before proceeding, notice that in the absence of a magnetosphere, i.e. ${\mathcal{P}}=0$, one recovers the standard gravitational frame-dragging effect of the Kerr black hole
$\Omega^{\text{K}}_{\infty}(r,\theta)\sim\frac{2aM}{r^{3}}$ (4.28)
such that for neutral particles sufficiently close to the black hole, the frame-dragging forces them to rotates with the hole.
The magnetic monopole dramatically changes this effect. When the magnetosphere is turned on, the large distance behavior is dominated by the electromagnetic interaction such that the azimuthal speed experienced by charged particle in our magnetized Kerr (MK) reads
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}\frac{(1-\cos{\theta})}{\sin^{2}{\theta}}$ (4.29)
We observer that for ${\mathcal{P}}>0$, the angular velocity has the opposite sign of $\kappa$, i.e. the magnetosphere forces the positive charges to have a retrograde motion while negatve charge have a prograde motion. For a collimated jet, i.e. for a small opening angle $\theta_{m}$, it reduces to
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{2\varepsilon r^{2}}\left(1+\frac{\theta_{m}^{2}}{4}\right)+{\mathcal{O}}(\theta_{m}^{4})$ (4.30)
Remarkably, it depends solely on the magnetic monopole charge $\mathcal{P}$}, the mass to charge ratio $\kappa$ and the energy of the particle $\varepsilon$. At the equator $\theta=\pi/2$, one finds
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}$ (4.31)
To conclude, the magnetosphere induces a new magnetic frame-dragging effect which completely dominates its usual gravitational frame dragging counterpart. Moreover, one notice that while the gravitational frame dragging effect decays as $1/r^{3}$ at large distance, the magnetic frame dragging decays only as $1/r^{2}$, showing that it remains effective on much larger distance away from the black hole.
Blueshift of charged particles and the zone of acceleration
We are now interested in characterizing what a Carter observer would measured when a charged particle crosses its trajectory. For that purpose, it will reveal useful to introduce the redshift $z$ by
$1+z=\frac{E_{\text{em}}}{E_{\text{obs}}},$ (4.32)
where $E_{\text{em}}$ is the energy of the particle when emitted and $E_{\text{obs}}$ the energy of the particle when observed, both energies being measured by the Carter observer.
The particle is blueshifted when $E_{\text{obs}}>E_{\text{em}}$ which corresponds to $z<0$. Let us denote $r_{\text{em}}$ the radius at which the particle is emitted and let us consider that the particle is observed at $r\rightarrow+\infty$. We recall that the Carter observer has a $4$-velocity given by $e_{0}$ [cf. Eq. (2.14a)]. The redshift factor (4.32) can thus be written as
$1+z=\frac{[u^{\mu}(e_{0})_{\mu}]_{\text{em}}}{[u^{\mu}(e_{0})_{\mu}]_{\infty}}.$ (4.33)
We read on the components (4.6) of $u$ with respect to the Carter tetrad that
$u^{\mu}(e_{0})_{\mu}=\frac{\varepsilon(r^{2}+a^{2})-a\ell}{\sqrt{\Sigma\Delta}}.$ (4.34)
This formula yields $u^{\mu}(e_{0})_{\mu}]_{\infty}=\varepsilon$. Hence we recover that the conserved quantity $\varepsilon$ is interpretable as the particle’s specific energy measured by an inertial observer at infinity and at rest with respect to the black hole (such as the asymptotic Carter observer). For the case $\ell=-\kappa{\mathcal{P}}$ that we considering [cf. Eq. (4.15)], we may write $\ell=a\beta$ [cf. Eq. (4.18)] and a direct computation then gives
$z=\frac{1}{\sqrt{\Sigma\Delta}}\left[r_{\text{em}}^{2}+a^{2}\left(1-\frac{\beta}{\varepsilon}\right)-\sqrt{(r_{\text{em}}^{2}+a^{2}\cos^{2}{\theta})(r_{\text{em}}^{2}-2Mr_{\text{em}}+a^{2})}.\right]$ (4.35)
The particle is blueshifted when the numerator is negative. This happens for sufficiently large values of $\beta/\varepsilon=-\kappa{\mathcal{P}}/(a\varepsilon)$ and $r_{\text{em}}$ sufficiently small, as shown in Figs. 34.
Figure 3: Redshift factor $z$ along the Northern rotation axis ($\theta=0$) as a function of the emission radius $r_{\rm em}$, for various values of the ratio $\beta/\varepsilon$ and for $a/M=0.1$ (top figures) and $a/M=0.5$ (bottom figures). Since $z$ is diverging for $r\to r_{+}$ (the horizon), the plots in the left figures start at $r=1.01\,r_{+}$.
Figure 4: Same as Fig. 3 but for $a/M=0.9$ (top figures) and $a/M=0.98$ (bottom figures).
First, fixing the spin of the black hole $a/M$, we observe that the extent of the region from which ejected particles appear blueshifted at infinity, i.e. $z<0$, grows with the ratio $\beta/\varepsilon$. For a slowly rotating black hole, i.e. $a=0.1M$ (top of Fig. 3), ejected particles are blueshifted only for very high values of the parameter $\beta/\varepsilon$. For higher value of the spin (bottom of Fig. 3 and Fig. 4), the required values of the parameter $\beta/\varepsilon$ for having a region of blueshift are much smaller. This behavior demonstrates the combined roles of both the spin and the magnetization of the black hole in accelerating charged particles for asymptotic observers.
Moreover, we observe on Figs. 34 that the region in which emitted particles are blueshifted admits a maximal radius, which corresponds to $z=0$. This maximal radius is always close to the horizon. For instance, for $a=0.1M$ and $\beta/\varepsilon=600$ or $a=0.5M$ and $\beta/\varepsilon=25$, the maximal radius is at $r_{\ast}\sim 5.8M>r_{+}$. The existence of this bounded region allowing charged particles to be blueshifted at infinity and its analytical description stand as the main results of this work. Within this specific model, this shows that accelerated particles of a given energy can only emerge from a finite region surrounding the black hole horizon.
Discussion
In this work, we have studied the ejection of charged particles in the magnetized Kerr black hole and the built-up of a relativistic jet near the poles. Contrary to most approaches, our investigation is fully analytic and provides close expressions for key quantities characterizing the jet (acceleration of the charged particles, polar equilibrium, frame-dragging effect, blueshift of the particle energy). Let us briefly summarized the main points of this analytic model.
The present model relies on the simplest Kerr magnetosphere which consists in the vacuum magnetic monopole configuration. As we reviewed in the Section 3.1, this exact solution is selected from the fundamental symmetries of the Kerr geometry and corresponds to the Penrose current (2.56) of the Kerr black hole. Therefore, it is intimately tied to the existence of a Killing-Yano tensor in this geometry. In turn, this exact solution of the source-free Maxwell equations in the Kerr geometry can also be recovered as the test field limit of the Kerr-Newman magnetosphere with a non-zero magnetic charge and a zero electric one. The key motivation for focusing on this Kerr magnetosphere model is that it belongs to the few exact solutions for which the motion of charged particles is fully separable. In Section 3, we have reviewed in detail the geometrical condition (3.30) for a magnetosphere to preserve the Carter constant which ultimately ensures the separability, and have shown that among the models of magnetosphere built from the Killing symmetries, only the Killing-Maxwell system, based on the principal tensor, and the Penrose current, based on the KY tensor, satisfy this key condition. The first two sections thus provide a general self-contained review on (i) the different models of Kerr magnetospheres based on the Killing symmetries and (ii) the necessary conditions for the electrogeodesic motion to be separable. While most of the results presented in these first two sections are known, we have also included new results. In particular, in Section 2.1.3, we have shown that besides the Wald construction, one can also use the commuting Killing vectors and the Hodge duality to build a new non-vacuum exact solution of the test Maxwell equations. This provides the solution (2.42), which describes a radially growing toroidal magnetic field surrounding the Kerr geometry. Whether this new solution can be useful at the vicinity of the hole remains to be studied. This concludes the summary of the first two sections intended to review the required material for our investigation of the motion of charged particles in the magnetized Kerr black hole. Let us now summarize the application to the jet launching mechanism.
Using the analytic solution of the electrogeodesic equation in the Kerr monopole vacuum magnetosphere, we have studied several key properties of the jet.
  • The form of the acceleration: We have presented the analytic expression for the $4$-velocity $u^{\mu}\partial_{\mu}$ [Eq. (4.6)] and 4-acceleration $a^{\mu}\partial_{\mu}$ [Eq. (4.7)] of the charged particles in the whole black hole exterior, which allowed us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, such that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of jet of charged particles around the poles. For ${\mathcal{P}}>0$ and negatively charged particles, $\kappa<0$, the acceleration is outwards and the particles are ejected from the North pole.
  • Polar equilibrium positions: We have analyzed the stable and unstable equilibrium positions of the polar motion depending on the energy and angular momentum $(\epsilon,\ell)$ of the charged particle. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$ where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the pole $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we have identified a new condition (4.22) for the (North) rotation axis to be a stable polar position: the energy of the particle must be either smaller or greater than a threshold depending on $\beta$, i.e (4.23). Instead, trajectories violating this condition admit a stable polar position at some $\theta_{\ast}\in\;]0,\pi/2[$.
  • Magnetic frame-dragging effect: We have studied the fate of the frame-dragging effect in the Kerr magnetic monopole. Using the expression of the azimuthal angular velocity, i.e. (4.27), we have shown that the magnetic monopole introduces a new magnetic frame-dragging which completely dominates the gravitational frame-dragging. Indeed, we show that the former decays as $1/r^{2}$ while the later decays as $1/r^{3}$, showing that the new magnetic frame-dragging impacts charged particles on larger scales, even in this test field approximation. Furthermore, in contrast to the gravitational frame-dragging, which forces particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their charge, i.e. on $\kappa$.
  • Geometry of the acceleration zone and asymptotic blueshift of charged particles: We have further characterized the condition under which charged particles ejected from the magnetosphere reach as asymptotic observer with a higher energy. By evaluating the redshift $z$ experienced by the particle [Eq. (4.35)] we have provided an analytical condition, i.e. $z>0$, for the particle to be blueshifted at large distance. Interestingly, this condition reveals the existence of a structure, through a radius which encodes the maximal extent of the region in which emitted particles escaping to infinity will be observed as blueshifted. The maximal radius of this region grows with the spin and the magnetization of the black hole.
  • Numerical values for the parameters: We have checked that our model is self-consistent for concrete realistic numerical values of the parameters. For an electron [^8] In terms of SI units, one has actually $\kappa=\sqrt{\mu_{0}/(4\pi)}\,q/m$, so that, for an electron, $\kappa_{\rm e}=-5.56\;10^{7}\;{\rm m}^{1/2}{\rm kg}^{-1/2}$, which yields $\kappa_{\rm e}=-2.04\;10^{21}$ in geometrical units., $\kappa=\kappa_{\rm e}=-2.04\;10^{21}$, while for a proton, $\kappa=\kappa_{\rm p}=1.11\;10^{18}$. These huge values reflect the large dominance of the electromagnetic interaction over the gravitational one for elementary particles. On the contrary, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$, in order for the electromagnetic field to not significantly modify the spacetime metric, which is assumed to be the Kerr metric; otherwise one would have to consider the Kerr-Newman metric. For values of $a\sim M$, this is equivalent to demand $|{\mathcal{P}}|/a\ll 1$. It follows that $\beta$ can take values of order one or even much higher. For instance, for electrons, for which $|\kappa|\sim 10^{21}$, $\beta$ can be as large as $10^{10}$ for $|{\mathcal{P}}|=10^{-10}M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
To our knowledge, the present model provides the simplest analytical study for the ejection of charged particles from the Kerr black hole (test) magnetosphere. It offers a first minimal model to study the jet launching within a vacuum Kerr magnetosphere based solely on the electrogeodesic motion, thus addressing one of the key challenge of field mentioned in the introduction. There are several open interesting directions suggested by our results.
The present investigation is only preliminar and only a subset of the electrogeodesics has been discussed, namely the one with $\ell=-\kappa{\mathcal{P}}$. It reveals a rich phenomenology, which needs to be studied in detail. A full classification of the different possible trajectories of charged particles will be discussed elsewhere. It would also be interesting to confront the present findings with numerical GR/PIC simulations of this magnetosphere. Finally, we point that the present model can serve as a background to perform a perturbative study of the magnetosphere and the ejection of particles.
Acknowledgments
J. BA is grateful to the University of Santiago de Chile (USACH) for its hospitality during his different visits where this project was worked out. J.BA also warmly thank Léa Jouvin for enjoyable and inspiring discussions on the galactic center and the $\gamma$-astronomy on the road to Tilcara. J. BA and I. EM are also grateful to Abdelkader Abdallah for motivating discussions in the first step of this work. I.EM acknowledges support from the grant ANID FONDECYT 11240206 and funding via the BASAL Centro de Excelencia en Astrofisica y Tecnologias Afines (CATA) grant PFB06/2007. E.G. acknowledges funding by l’Agence Nationale de la Recherche, projects StronG ANR-22-CE31-0015-01 and Einstein PPF ANR-23-CE40-0010-02.
Appendix A SageMath notebooks
Some computations performed in this article have been performed by means the free Python-based mathematical software system SageMath [80, 81]. The relevant Jupyter notebooks are publicly available, from the links below:
  1. 1.
    Killing and Killing-Yano magnetospheres arround a Kerr black hole:
    https://nbviewer.org/url/relativite.obspm.fr/notebooks/KY_magnetosphere.ipynb
  2. 2.
    The Kerr monopole magnetosphere and its electrogeodesics:
    https://nbviewer.org/url/relativite.obspm.fr/notebooks/Kerr_monopole.ipynb